# UniVST: A Unified Framework for Training-free Localized Video Style Transfer

### v3

@@ 6 Multi-Step Video Style Transfer
Apart from the cases presented in the experimental section, UniVST showcases its versatility by enabling multi-step video style transfer, which allows for more customized transformations and creative potential of video editing. For instance, users can start by stylizing the background of the video using a mask to isolate it from the rest of the scene. After completing this step, they can invert the mask to focus on the previously excluded foreground and apply a separate style transfer to it. This sequential process not only provides greater control but also enables creative layering of styles. As illustrated in Fig. 15, employing distinct style images for the background and foreground in these two stages produces a visually striking and dynamic effect, enhancing the artistic flexibility of the method.
@@ 7 Limitation and Future Work
While our UniVST framework shows a clear advantage in localized video style transformation, it has limitations: (1) The accuracy of mask propagation and video content fidelity may be affected by inaccuracies in the inversion process. (2) The style transfer mechanism, guided solely by prior knowledge of LDM [29], may struggle with certain style images, especially those with nuanced color distributions, leading to suboptimal performance. (3) The sliding window smoothing technique, based on optical flow, may be less effective with fast-moving objects, resulting in unsatisfactory outcomes. Future research could focus on developing improved transfer techniques that may enhance visual effects and better manage scenes with rapid motion.
@@ 8 Conclusion
This paper presents UniVST, a unified framework for localized video style transfer that features a training-free approach, marking a significant advancement over traditional methods that require full video style transfer. Our contributions include: (1) A point-matching mask propagation strategy that eliminates the need for tracking models, enabling streamlined style transfer to specific video objects. (2) A training-free AdaIN-guided video style transfer mechanism that operates at both the latent and attention levels, ensuring a balance between content fidelity and style richness. (3) A sliding-window consistent smoothing scheme that uses optical flow to refine noise and update the latent space, improving temporal consistency and reducing artifacts. Extensive experiments show that UniVST outperforms existing methods in both qualitative and quantitative assessments, preserving the primary object’s style while ensuring temporal consistency and detail preservation. Futhermore, it can be extended to a wider range of customized models.

### v4

@@ 5.7 Multi-Step Video Style Transfer
Apart from the cases presented in the experimental section, UniVST showcases its versatility by enabling multi-step video style transfer, which allows for more customized transformations and the creative potential of video editing. For instance, users can start by stylizing the background of the video using a mask to isolate it from the rest of the scene. After completing this step, they can invert the mask to focus on the previously excluded foreground and apply a separate style transfer to it. This sequential process not only provides greater control but also enables creative layering of styles. As illustrated in Fig. 17, employing distinct style images for the background and foreground in these two stages produces a visually striking and dynamic effect, enhancing the artistic flexibility of the method.
@@ 6 Limitation and Future Work
While our UniVST framework shows a clear advantage in localized video style transformation, it also has limitations: (1) The accuracy of mask propagation and video content fidelity may be significantly affected by inaccuracies in the inversion process. (2) The style transfer mechanism, guided solely by prior knowledge of LDM [32], may struggle with certain style images, especially those with nuanced color distributions, leading to suboptimal performance. (3) The sliding-window smoothing technique, based on optical flow, can introduce additional inference time. Future research could explore improved stylization techniques to enhance visual effects and reduce inference overhead from optical flow prediction.
@@ 7 Conclusion
This paper presents UniVST, a unified framework for localized video style transfer without additional training, marking a significant advancement over traditional methods that require full video style transfer. Our contributions include: (1) A point-matching mask propagation strategy that eliminates the need for tracking models, enabling streamlined style transfer to specific video objects. (2) A training-free AdaIN-guided localized video stylization mechanism that operates at both the latent and attention levels, ensuring a balance between content fidelity and style richness. (3) A sliding-window consistent smoothing scheme that uses optical flow to refine noise and update the latent space, improving temporal consistency and reducing artifacts. Extensive experiments show that our UniVST outperforms existing methods in both qualitative and quantitative evaluations, effectively preserving the style of the primary object while ensuring temporal consistency. Furthermore, it can be extended to a broader range of customized models.